le tine direct materials used during the year.) E2-26A Compute Cost of Goods Manufactured and Cost of Goods Soled (Learning Objective 5) Compute the Cost of Goods Manufactured and Cost of Goods Sold for West Nautical Company for the most recent year using the amounts described next. Assume that the Raw Materials Inventory contains only direct materials Beginning End End of Year of Year of Year Raw materials inventory. Work in process inventory Finished goods inventory Purchases of direct materials.... Direct labor Indirect labor. $23,000 $25,000 Insurance on plant. Depreciation-plant building and equipment Repairs and maintenance-plant $13,400 S 3,700 . 77,000 $28,500 $35,000 $31,000 $20,000 $22,000 $74,000 Marketing expenses $86,000 General and administrative expenses.... $42,000 E2-27A Continues E2-26A: Prepare income statement (Learning Objective 5) Prepare the income statement for West Nautical Company in E2-26A for the most recent year. Assume that the company sold 34,000 units of its product at a price of $12 each during the year.
Answer #1
E2-26A
Calculation of Cost of Goods Manufactured
For Current Year
Beginning work in process inventory 35000
Plus: Manufacturing costs incurred
Direct materials used 72000 =23000+74000-25000
Direct labor 86000
Manufacturing overhead 70600 =42000+11500+13400+3700
Total manufacturing costs to account for 263600
Less: Ending work in process inventory 31000
Cost of goods manufactured (CGM) 232600
Calculation of Cost of Goods Sold
For Current Year
Beginning finished goods inventory 20000
Plus: Cost of goods manufactured 232600
Cost of goods available for sale 252600
Less: Ending finished goods inventory 22000
Cost of goods sold 230600
E2-27A
Income Statement
For Current Year
Sales revenues 408000 =34000*12
Less: Cost of goods sold 230600
Gross profit 177400
Less: Operating expenses
Marketing expenses 77000
General and administrative expenses 28500 105500
Operating income 71900